Lucia's aunt is now three times as old as Lucia. Fifteen years ago, her aunt was five times as old as Lucia. How old are Lucia and her aunt?

Answer by ankor@dixie-net.com(12704)

Lucia's aunt is now three times as old as Lucia. Fifteen years ago, her aunt was five times as old as Lucia. How old are Lucia and her aunt?
:
Let x = Lucia's age now
Then
3x = Auntie's age now
:
15 years ago they would be:
(x-15) = Lucia
and
(3x-15) = Auntie
:
It says,"15 yrs ago aunt was five times as old as Lucia."
(3x-15) = 5(x-15)
3x - 15 = 5x - 75
3x - 5x = -75 + 15
-2x = -60
x = -60/-2
x = +30 yrs is Lucia's age now
Then
3(30) = 90 yrs is Auntie's age now.
:
Check solutions by looking at their ages 15 years ago
Lucia = 15 years
Auntie = 75 years which is 5 times 15 years